# LD-MoLE: Learnable Dynamic Routing for Mixture of LoRA Experts

### v1

@@ Abstract
Recent studies have shown that combining parameter-efficient fine-tuning (PEFT) with mixture-of-experts (MoE) is an effective strategy for adapting large language models (LLMs) to the downstream tasks. However, most existing approaches rely on conventional TopK routing, which requires careful hyperparameter tuning and assigns a fixed number of experts to each token. In this work, we propose LD-MoLE, a Learnable Dynamic routing mechanism for Mixture of LoRA Experts that enables adaptive, token-dependent, and layer-wise expert allocation. Our method replaces the non-differentiable TopK selection with a differentiable routing function and a closed-form solution. Moreover, our design allows the model to adaptively determine the number of experts to activate for each token at different layers. In addition, we introduce an analytical sparsity control objective to regularize the number of activated experts. Extensive experiments on the Qwen3-1.7B and Llama-3.2-3B models show that LD-MoLE achieves the highest average scores compared to state-of-the-art baselines, across a diverse set of benchmarks. Our method not only achieves superior performance, but also demonstrates the ability to learn token-dependent and layer-wise expert allocation.
@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive capabilities across a wide range of natural language processing (NLP) tasks. However, their growing size requires significant computational resources for full-parameter fine-tuning. To address this, Parameter-Efficient Fine-tuning (PEFT) methods, such as Adapter-tuning (Houlsby et al., 2019) and LoRA (Hu et al., 2021), have emerged as crucial techniques for reducing training costs.
@@
Recently, the Mixture-of-Experts (MoE) design (Jacobs et al., 1991; Shazeer et al., 2017) has been successfully integrated into transformer feed-forward networks during LLMs pretraining (Dai et al., 2024; Yang et al., 2025), demonstrating that MoE can reduce computational cost while maintaining strong performance. This has inspired a promising direction for PEFT, leading to the Mixture of LoRA Experts (MoLE) framework (Wu et al., 2024; Dou et al., 2024; Zadouri et al., 2023). MoLE utilizes multiple LoRAs as experts, providing a scalable and efficient alternative to relying on a single LoRA – where high-rank configurations risk overfitting and increased compute cost (Zhang et al., 2023), while low-rank ones often underperform (Liao et al., 2025; Gao et al., 2024).
@@
Despite substantial advances, many recent MoE variants remain constrained by rigid routing strategies. A prominent example is MoLA (Gao et al., 2024), which relies on conventional TopK routing. This approach forces every token to consult a fixed number of experts, introducing a manually tuned hyperparameter that prevents adaptive allocation of resources based on token complexity. In addition, the discrete and non-differentiable nature of the TopK operator hinders end-to-end optimization, ultimately limiting both performance and scalability (Shazeer et al., 2017; Zoph et al., 2022; Wang et al., 2025). Recent efforts such as ReMoE (Wang et al., 2025) attempt to bypass this bottleneck by replacing TopK with a ReLU-based router, but this dynamic scheme can suffer from instability, as some tokens may be routed to no experts at all, degrading overall performance. Taken together, these limitations highlight a central challenge: Can we design a routing mechanism that adaptively learns to allocate experts in a stable and differentiable way?
@@
In this work, we propose LD-MoLE (see Figure 1), a Learnable and Dynamic routing method to adaptively control LoRA experts allocation. We adopt Sparsegen (Laha et al., 2018) as the projection onto the probability simplex and propose a dynamic routing mechanism and the corresponding training pipeline that has the following benefits: (1) the closed-form formulation with Sparsegen to decide routing probability ensures differentiability and guarantees that every token is assigned to at least one expert; (2) the routing admits a well-defined subgradient; (3) the derivative of the routing is upper-bounded, facilitating stable optimization; and (4) the routing design supports sparse yet controllable allocations. Building on this foundation, we introduce a lightweight, shared multi-layer perceptron (MLP) that predicts a token-specific sparsity parameter $\lambda$, governing expert selection. In addition, we formulate a sparsity control objective derived from Sparsegen’s analytical solution, enabling direct regularization over the number of activated experts.
@@
We conduct extensive experiments to validate the effectiveness of LD-MoLE. Specifically, we adopt Llama-3.2-3B and Qwen3-1.7B as base LLMs and fine-tune them on a wide range of instruction-tuning and sequence classification benchmarks. LD-MoLE achieves the best performance across these benchmarks, outperforming prominent baselines that follow different routing strategies –MoLA (Gao et al., 2024) with conventional TopK routing and ReMoE (Wang et al., 2025) with ReLU-based dynamic routing.These results indicate that our learnable routing mechanism yields consistent improvements across tasks and architectures. Moreover, we show that our sparsity control loss effectively reduces the number of activated experts without compromising performance.
@@
    We conduct comprehensive experiments on Llama-3.2-3B and Qwen3-1.7B, including ablation studies and detailed analyses, to demonstrate the effectiveness of LD-MoLE and to elucidate the mechanisms behind its improvements over TopK and ReLU-based routing.
@@ 2 Related Work
MoE was first introduced in the 1990s (Jacobs et al., 1991) and later applied to large-scale neural networks (Shazeer et al., 2017) to efficiently scale up model capacity. Landmark models like Google’s GShard (Lepikhin et al., 2020) implement sparse MoE frameworks with Top2 expert routing, while Switch Transformer (Fedus et al., 2022) simplifies this to a single expert per token to reduce overhead. Today, MoE has been widely adopted in several well-known large language models, including GLaM (Du et al., 2022), Mixtral-8x7B (Jiang et al., 2024), DeepSeekMoE (Dai et al., 2024), and Qwen3 (Yang et al., 2025).
@@
Various routing strategies have been proposed for expert selection. The most common is TopK routing (Shazeer et al., 2017), where each token selects a fixed number of experts. Alternative designs, such as expert-choice routing (Zhou et al., 2022) reverse this perspective by allowing experts to select tokens. Beyond fixed-$k$ approaches, several methods aim to enable dynamic routing. For instance, TopP routing (Huang et al., 2024) selects experts until a cumulative probability threshold is reached, while DYNMOE (Guo et al., 2025) introduces Top-Any Gating to eliminate the need for tuning $k$. Soft MoE (Puigcerver et al., 2024) merges tokens and assigns them to experts as linear combinations, and Lory (Zhong et al., 2024) proposes a fully differentiable routing mechanism but underperforms TopK routing. Closest to our work, ReMoE (Wang et al., 2025) employs ReLU-based routing for differentiable and dynamic selection.
@@
Combining multiple LoRA modules (Hu et al., 2021) with MoE structure has led to the Mixture of LoRA Experts framework (Wu et al., 2024). Several variants have since been proposed: LoRAMoE (Dou et al., 2024) introduces MoE-style plugins to enhance downstream performance while mitigating knowledge forgetting; HMoRA (Liao et al., 2025) employs a hybrid scheme that hierarchically integrates token-level and task-level routing. MixLoRA (Li et al., 2024) builds a resource-efficient sparse MoE from LoRA modules. Other works, such as MoLA (Gao et al., 2024) and AlphaLoRA (Qing et al., 2024), analyze expert allocation patterns across layers. In this work, we introduce LD-MoLE, which integrates the fully differentiable Sparsegen formulation with a learned MLP to predict $\lambda$, enabling end-to-end dynamic expert routing.
@@ 3.1 The LD-MoLE Architecture
which yields a sparse probability vector with at most $k$ nonzero entries. Note that the selection operator $\mathcal{S}_{k}(\cdot)$ introduces a jump discontinuity at the $k$-th largest value, making the routing function non-differentiable.
Despite the success of conventional TopK routing with softmax operation in improving training and inference efficiency, two limitations persist (Guo et al., 2025; Wang et al., 2025): (1) TopK routing is non-differentiable during the learning process. (2) The value of $k$ requires carefully tuned to optimize model performance and would be fixed throughout the training process  (Guo et al., 2025).
@@
To address the aforementioned limitations, we propose a learnable dynamic routing mechanism based on Sparsegen (Laha et al., 2018), which is a projection onto the probability simplex that generates sparse outputs via a closed-form and fully differentiable solution. Given the score vector $\bm{u}$} in Eq. 1, the routing function adaptively determines the effective number of activated experts by solving a closed-form transformation where $\lambda$ is a sparsity scalar:
@@
Proposition 1 (Closed-form Sparsegen routing: Proposition 0.1 in (Laha et al., 2018)).
@@
This result follows from solving the Sparsegen projection problem, which minimizes a strongly convex objective subject to simplex constraints (Laha et al., 2018). ∎
@@ 3.2 Training Loss
In this work, we adopt the standard cross-entropy loss for the Language Model (LM) in both next-token prediction and sequence classification tasks (Xue et al., 2024; Liao et al., 2025; Wu et al., 2024; Dou et al., 2024). Formally, this could be expressed as
@@
To further stabilize the training process, we incorporate the conventional load-balancing loss (Fedus et al., 2022; Yang et al., 2025; Dai et al., 2024), which mitigates the risk of routing collapse (Shazeer et al., 2017). Such collapse can also arise in LoRA-augmented expert settings during fine-tuning, where only a few experts dominate the token assignments. Additionally, we introduce a sparsity loss that leverages the closed-form nature of our routing to directly regulate the sparsity level. In the following, we present the mathematical formulation of both the load-balancing loss and the proposed sparsity loss in detail.
@@ 4.1 Experiment Setup
Benchmarks and Metrics: We evaluate the overall accuracy of our method against several baselines across a range of downstream tasks. Specifically, we test on standard NLP benchmarks, including instruction-finetuning datasets such as ARC-Challenge, ARC-Easy (Clark et al., 2018), OpenBookQA (Mihaylov et al., 2018), CommonsenseQA (Talmor et al., 2019), SWAG (Zellers et al., 2018), HellaSWAG (Zellers et al., 2019), as well as sequence classification tasks from GLUE: CoLA, and RTE (Wang et al., 2019). For all benchmarks, we use standard accuracy as the evaluation metric. Please refer to Appendix E for more detail of the dataset and setup.
Base Model and Baselines: We test baseline approaches on different open-source LLMs, including Llama-3.2-3B (Dubey et al., 2024) and Qwen3-1.7B (Yang et al., 2025). We compare our method primarily against MoLA (Gao et al., 2024), a TopK routing strategy within the MoLE framework, denoted as MoLA(8888). We also evaluate its proposed variant MoLA(2468), which assigns fewer experts to lower layers and progressively increases the allocation toward higher layers, reportedly yielding consistently better performance. In addition, we include ReMoLE, which adapts the ReLU-based routing from ReMoE (Wang et al., 2025) to the LoRA experts setting. Simliar to L2D-MoLE, ReMoLE supports both dynamic and differentiable routing.
Implementation: For our method, training is conducted on 4 NVIDIA H200 GPUs with a batch size of 16 for 10 epochs, with the learning rate of 0.0001 decayed by a factor of 0.1 at epochs 6 and 8. We set the number of LoRA experts to 8, with rank 8 and scaling factor 16, and apply a dropout rate of 0.1. Across all methods, we pair the training with the load-balancing loss and follow the settings described in the original baseline papers. For MoLA, we choose the top 2 expert selections follows the original settings (Gao et al., 2024). For ReMoLE, we employ the load-balancing objective function introduced in ReMoE(Wang et al., 2025) with exact the same coefficients. More details of the method and experiment training setting are provided in Appendix D.
@@ 4.2 Overall performance
The overall performance of our proposed LD-MoLE is summarized in Table 1. Across all tested configurations, LD-MoLE achieves the highest average scores on both the Llama-3.2-3B and Qwen3-1.7B models, demonstrating the consistent benefits of its learned dynamic routing. For this comparison, we set $\alpha=1.0$ and disable the sparsity loss (i.e., $\beta=0$), as defined in Eq. 15. A detailed analysis of sparsity control is deferred to Section 4.4.
@@
In particular, LD-MOLE outperforms both fixed and dynamic routing baselines. Compared to the fixed TopK routing of MoLA, our method excels on reasoning-heavy benchmarks, achieving average cross-model gains of over +3.5% on ARC-E, SWAG, and HellaSWAG. On OpenBookQA, it achieves an average improvement of about +1.2%, and on CommonsenseQA, it surpasses MoLA by more than +2.0%. While MoLA attains slightly better results on certain sequence classification tasks such as RTE, LD-MoLE consistently delivers higher overall averages, underscoring the effectiveness of learnable dynamic routing across diverse task types. Compared to ReMoLE, LD-MoLE achieves higher overall averages, including +0.5% on Llama3-2.3B and +0.6% on Qwen3-1.7B. Notably, ReMoLE exhibits large performance drops on CoLA with Llama3-2.3B and RTE with Qwen3-1.7B, whereas LD-MoLE maintains stable effectiveness across benchmarks.
@@
We observe that dynamic routing methods generally perform better on instruction fine-tuning tasks, while fixed routing approaches show slight advantages in certain sequence classification tasks. A possible explanation is that classification tasks often benefit from more uniform expert usage, where fixed routing ensures stable allocation. Interestingly, the pruned variant MoLA(2468) outperforms the standard MoLA(8888), suggesting that many experts in the fixed routing setup are underutilized, introducing redundancy as also noted in their work  (Gao et al., 2024). In contrast, dynamic routing adapts expert selection on a token-by-token basis, which benefits complex reasoning and instruction-following tasks but may introduce variability that is less advantageous for shorter classification settings. Overall, LD-MoLE provides a stronger balance between parameter efficiency and performance, adapting better effectiveness across diverse tasks.
@@
Method Model TP ARC-C ARC-E Open Comm SWAG HellaSWAG CoLA RTE Avg
MoLA(8888) Llama3.2-3B 3.11 % 71.57 83.51 81.00 79.77 83.56 87.47 85.81 90.61 82.91
MoLA(2468) Llama3.2-3B 1.80 % 71.91 83.86 83.60 80.02 83.96 87.31 86.00 89.53 83.27
ReMoLE Llama3.2-3B 3.11 % 75.25 89.30 83.40 79.52 90.45 93.44 83.95 89.46 85.59
LD-MoLE Llama3.2-3B 3.28 % 74.58 89.47 84.00 81.42 91.37 93.60 86.02 88.38 86.10
MoLA(8888) Qwen3-1.7B 4.12 % 76.59 88.60 82.40 76.49 84.11 83.35 83.89 86.64 82.75
MoLA(2468) Qwen3-1.7B 2.39 % 76.92 88.42 83.00 75.84 84.17 87.09 83.60 84.48 82.94
ReMoLE Qwen3-1.7B 4.12 % 79.60 91.75 84.80 79.44 86.37 88.00 82.12 83.74 84.47
LD-MoLE Qwen3-1.7B 4.23 % 78.67 92.11 85.00 79.30 86.72 88.71 82.61 87.72 85.10

### v2

@@ Abstract
Recent studies have shown that combining parameter-efficient fine-tuning (PEFT) with mixture-of-experts (MoE) is an effective strategy for adapting large language models (LLMs) to the downstream tasks. However, most existing approaches rely on conventional TopK routing, which requires careful hyperparameter tuning and assigns a fixed number of experts to each token. In this work, we propose LD-MoLE, a Learnable Dynamic routing mechanism for Mixture of LoRA Experts that enables adaptive, token-dependent, and layer-wise expert allocation. Our method replaces the non-differentiable TopK selection with a differentiable routing function and a closed-form solution. Moreover, our design allows the model to adaptively determine the number of experts to activate for each token at different layers. In addition, we introduce an analytical sparsity control objective to regularize the number of activated experts. Extensive experiments on the Qwen3-1.7B and Llama-3.2-3B models show that LD-MoLE achieves the highest average scores compared to state-of-the-art baselines, across a diverse set of benchmarks. Our method not only achieves superior performance, but also demonstrates the ability to learn token-dependent and layer-wise expert allocation. The code is available at: https://github.com/eshentw/LD-MoLE.
@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive capabilities across a wide range of natural language processing (NLP) tasks. However, their growing size requires significant computational resources for full-parameter fine-tuning. To address this, Parameter-Efficient Fine-tuning (PEFT) methods, such as Adapter-tuning (houlsbyParameterEfficientTransferLearning2019a) and LoRA (LoRA), have emerged as crucial techniques for reducing training costs.
@@
Recently, the Mixture-of-Experts (MoE) design (originalMoE; shazeerOutrageouslyLargeNeural2017a) has been successfully integrated into transformer feed-forward networks during LLMs pretraining (daiDeepSeekMoEUltimateExpert2024; yangQwen3TechnicalReport2025), demonstrating that MoE can reduce computational cost while maintaining strong performance. This has inspired a promising direction for PEFT, leading to the Mixture of LoRA Experts (MoLE) framework (MoLE; douLoRAMoEAlleviateWorld2024; zadouri2023pushingmixtureexpertslimit). MoLE utilizes multiple LoRAs as experts, providing a scalable and efficient alternative to relying on a single LoRA – where high-rank configurations risk overfitting and increased compute cost (zhang2023adaloraadaptivebudgetallocation), while low-rank ones often underperform (liaoHMORAMAKINGLLMS2025; gaoHigherLayersNeed2024).
@@
Despite substantial advances, many recent MoE variants remain constrained by rigid routing strategies. A prominent example is MoLA (gaoHigherLayersNeed2024), which relies on conventional TopK routing. This approach forces every token to consult a fixed number of experts, introducing a manually tuned hyperparameter that prevents adaptive allocation of resources based on token complexity. In addition, the discrete and non-differentiable nature of the TopK operator hinders end-to-end optimization, ultimately limiting both performance and scalability (shazeerOutrageouslyLargeNeural2017a; ST-MoE; ReLUMoE). Recent efforts such as ReMoE (ReLUMoE) attempt to bypass this bottleneck by replacing TopK with a ReLU-based router, but this dynamic scheme can suffer from instability, as some tokens may be routed to no experts at all, degrading overall performance. Taken together, these limitations highlight a central challenge: Can we design a routing mechanism that adaptively learns to allocate experts in a stable and differentiable way?
@@
In this work, we propose LD-MoLE (see Figure 1), a Learnable and Dynamic routing method to adaptively control LoRA experts allocation. We adopt Sparsegen (sparsegen) as the projection onto the probability simplex and propose a dynamic routing mechanism and the corresponding training pipeline that has the following benefits: (1) the closed-form formulation with Sparsegen to decide routing probability ensures differentiability and guarantees that every token is assigned to at least one expert; (2) the routing admits a well-defined subgradient; (3) the derivative of the routing is upper-bounded, facilitating stable optimization; and (4) the routing design supports sparse yet controllable allocations. Building on this foundation, we introduce a lightweight, shared multi-layer perceptron (MLP) that predicts a token-specific sparsity parameter $\lambda$, governing expert selection. In addition, we formulate a sparsity control objective derived from Sparsegen’s analytical solution, enabling direct regularization over the number of activated experts.
@@
We conduct extensive experiments to validate the effectiveness of LD-MoLE. Specifically, we adopt Llama-3.2-3B and Qwen3-1.7B as base LLMs and fine-tune them on a wide range of instruction-tuning and sequence classification benchmarks. LD-MoLE achieves the best performance across these benchmarks, outperforming prominent baselines that follow different routing strategies –MoLA (gaoHigherLayersNeed2024) with conventional TopK routing and ReMoE (ReLUMoE) with ReLU-based dynamic routing.These results indicate that our learnable routing mechanism yields consistent improvements across tasks and architectures. Moreover, we show that our sparsity control loss effectively reduces the number of activated experts without compromising performance.
@@
    We conduct comprehensive experiments on Llama-3.2-3B, Llama-3.2-8b and Qwen3-1.7B, including ablation studies and detailed analyses, to demonstrate the effectiveness of LD-MoLE and to elucidate the mechanisms behind its improvements over TopK and ReLU-based routing.
@@ 2 Related Work
MoE was first introduced in the 1990s (originalMoE) and later applied to large-scale neural networks (shazeerOutrageouslyLargeNeural2017a) to efficiently scale up model capacity. Landmark models like Google’s GShard (GShard) implement sparse MoE frameworks with Top2 expert routing, while Switch Transformer (fedus2022switchtransformersscalingtrillion) simplifies this to a single expert per token to reduce overhead. Today, MoE has been widely adopted in several well-known large language models, including GLaM (du2022glam), Mixtral-8x7B (jiang2024mixtralexperts), DeepSeekMoE (daiDeepSeekMoEUltimateExpert2024), Qwen3 (yangQwen3TechnicalReport2025) and LongCat-Flash (meituanlongcatteam2025longcatflashtechnicalreport).
@@
Various routing strategies have been proposed for expert selection. The most common is TopK routing (shazeerOutrageouslyLargeNeural2017a), where each token selects a fixed number of experts. There are also several works that discuss variants of TopK. AdaMoE (zeng2024adamoetokenadaptiveroutingnull) uses conventional TopK routing with k larger than in vanilla MoE but achieves token-adaptive expert selection by incorporating null experts, which are defined as an empty operation. Ada-K Routing (yue2024adakroutingboostingefficiency) introduces an allocator and then obtains $k^{*}$ for customized expert resource allocation instead of fixed TopK through a non-differentiable sampling operation with a RL-based optimization framework. Alternative designs, such as expert-choice routing (zhou2022mixtureofexpertsexpertchoicerouting) reverse this perspective by allowing experts to select tokens. Beyond fixed-$k$ approaches, several methods aim to enable dynamic routing. For instance, TopP routing (huang2024hardertasksneedexperts) selects experts until a cumulative probability threshold is reached, while DYNMOE (guoDynamicMixtureExperts2025) introduces Top-Any Gating to eliminate the need for tuning $k$. Soft MoE (puigcerverSparseSoftMixtures2024a) merges tokens and assigns them to experts as linear combinations, and Lory (zhongLoryFullyDifferentiable2024a) proposes a fully differentiable routing mechanism but underperforms TopK routing. Closest to our work, ReMoE (ReLUMoE) employs ReLU-based routing for differentiable and dynamic selection.
@@
Combining multiple LoRA modules (LoRA) with MoE structure has led to the Mixture of LoRA Experts framework (MoLE). Several variants have since been proposed: LoRAMoE (douLoRAMoEAlleviateWorld2024) introduces MoE-style plugins to enhance downstream performance while mitigating knowledge forgetting; HMoRA (liaoHMORAMAKINGLLMS2025) employs a hybrid scheme that hierarchically integrates token-level and task-level routing. MixLoRA (liMixLoRAEnhancingLarge2024a) builds a resource-efficient sparse MoE from LoRA modules. Other works, such as MoLA (gaoHigherLayersNeed2024) and AlphaLoRA (AlphaLoRA), analyze expert allocation patterns across layers. In this work, we introduce LD-MoLE, which integrates the fully differentiable Sparsegen formulation with a learned MLP to predict $\lambda$, enabling end-to-end dynamic expert routing.
@@ 3.1 The LD-MoLE Architecture
which yields a sparse probability vector with at most $k$ nonzero entries. Note that the selection operator $\mathcal{S}_{k}(\cdot)$ introduces a jump discontinuity at the $k$-th largest value. Consequently, an arbitrarily small perturbation of the router scores can change the selected set and induce an abrupt change in the gradient, rendering the routing function non-differentiable at these boundaries.
Despite the success of conventional TopK routing with softmax operation in improving training and inference efficiency, two limitations persist (guoDynamicMixtureExperts2025; ReLUMoE): (1) TopK routing is non-differentiable during the learning process. (2) The value of $k$ requires carefully tuned to optimize model performance and would be fixed throughout the training process  (guoDynamicMixtureExperts2025).
In contrast, Sparsegen (sparsegen) produces sparse routing weights via a closed-form projection, avoiding discrete TopK selection and yielding well-defined gradients that better align optimization with the routing behavior.
@@
To address the aforementioned limitations, we propose a learnable dynamic routing mechanism based on Sparsegen (sparsegen), which is a projection onto the probability simplex that generates sparse outputs via a closed-form and fully differentiable solution. Given the score vector $\bm{u}$} in Eq. 1, the routing function adaptively determines the effective number of activated experts by solving a closed-form transformation where $\lambda$ is a sparsity scalar:
@@
Proposition 1 (Closed-form Sparsegen routing: Proposition 0.1 in (sparsegen)).
@@
This result follows from solving the Sparsegen projection problem, which minimizes a strongly convex objective subject to simplex constraints (sparsegen). ∎
@@ 3.2 Training Loss
In this work, we adopt the standard cross-entropy loss for the Language Model (LM) in both next-token prediction and sequence classification tasks (OpenMoE; liaoHMORAMAKINGLLMS2025; MoLE; douLoRAMoEAlleviateWorld2024). Formally, this could be expressed as
@@
To further stabilize the training process, we incorporate the conventional load-balancing loss (fedus2022switchtransformersscalingtrillion; yangQwen3TechnicalReport2025; daiDeepSeekMoEUltimateExpert2024), which mitigates the risk of routing collapse (shazeerOutrageouslyLargeNeural2017a). Such collapse can also arise in LoRA-augmented expert settings during fine-tuning, where only a few experts dominate the token assignments. Additionally, we introduce a sparsity loss that leverages the closed-form nature of our routing to directly regulate the sparsity level. In the following, we present the mathematical formulation of both the load-balancing loss and the proposed sparsity loss in detail.
@@ 4.1 Experiment Setup
Benchmarks and Metrics: We evaluate the overall accuracy of our method against several baselines across a range of downstream tasks. Specifically, we test on standard NLP benchmarks, including instruction-finetuning datasets such as MMLU-Pro (wang2024mmluprorobustchallengingmultitask), ARC-Challenge, ARC-Easy (arc), OpenBookQA (openbookqa), CommonsenseQA (csqa), SWAG (swag), HellaSWAG (hellaswag), as well as sequence classification tasks from GLUE: CoLA, and RTE (glue). For all benchmarks, we use standard accuracy as the evaluation metric. Please refer to Appendix E for more detail of the dataset and setup.
Base Model and Baselines: We test baseline approaches on different open-source LLMs, including Llama-3.2-3B, Llama-3.2-8b (dubey2024llama) and Qwen3-1.7B (yangQwen3TechnicalReport2025). We compare our method primarily against MoLA (gaoHigherLayersNeed2024), a TopK routing strategy within the MoLE framework, denoted as MoLA(8888). We also evaluate its proposed variant MoLA(2468), which assigns fewer experts to lower layers and progressively increases the allocation toward higher layers, reportedly yielding consistently better performance. In addition, we include ReMoLE, which adapts the ReLU-based routing from ReMoE (ReLUMoE) to the LoRA experts setting. Simliar to L2D-MoLE, ReMoLE supports both dynamic and differentiable routing.
Implementation: For our method, training is conducted on 4 NVIDIA H200 GPUs with a batch size of 16 for 10 epochs, with the learning rate of 0.0001 decayed by a factor of 0.1 at epochs 6 and 8. For Llama-3.2-8b, we run the experiment for only run 3 epochs as we observe a convergence at the early stage of training. We set the number of LoRA experts to 8, with rank 8 and scaling factor 16, and apply a dropout rate of 0.1. Across all methods, we pair the training with the load-balancing loss and follow the settings described in the original baseline papers. For MoLA, we choose the top 2 expert selections follows the original settings (gaoHigherLayersNeed2024). For ReMoLE, we employ the load-balancing objective function introduced in ReMoE(ReLUMoE) with exact the same coefficients. More details of the method and experiment training setting are provided in Appendix D.
@@ 4.2 Overall performance
The overall performance of our proposed LD-MoLE is summarized in Table 1. Across all tested configurations, LD-MoLE achieves the highest average scores on both the Llama-3.2-3B, Llama-3.2-8b and Qwen3-1.7B models, demonstrating the consistent benefits of its learned dynamic routing. For this comparison, we set $\alpha=1.0$ and disable the sparsity loss (i.e., $\beta=0$), as defined in Eq. 15. A detailed analysis of sparsity control is deferred to Section 4.4.
@@
In particular, LD-MoLE outperforms both fixed and dynamic routing baselines. Compared to the fixed TopK routing of MoLA, our method excels on reasoning-heavy benchmarks, achieving average cross-model gains of over +3.5% on ARC-E, SWAG, and HellaSWAG. For MMLU-Pro, LD-MoLE outperform the baselines in all testing models. On OpenBookQA, it achieves an average improvement of about +1.2%, and on CommonsenseQA, it surpasses MoLA by more than +2.0%. While MoLA attains slightly better results on certain sequence classification tasks such as RTE, LD-MoLE consistently delivers higher overall averages, underscoring the effectiveness of learnable dynamic routing across diverse task types. Compared to ReMoLE, LD-MoLE achieves higher overall averages, including +0.5% on Llama3-2.3B and +0.6% on Qwen3-1.7B. Notably, ReMoLE exhibits large performance drops on CoLA with Llama3-2.3B and RTE with Qwen3-1.7B, whereas LD-MoLE maintains stable effectiveness across benchmarks.
@@
We observe that dynamic routing methods generally perform better on instruction fine-tuning tasks, while fixed routing approaches show slight advantages in certain sequence classification tasks. A possible explanation is that classification tasks often benefit from more uniform expert usage, where fixed routing ensures stable allocation. Interestingly, the pruned variant MoLA(2468) outperforms the standard MoLA(8888), suggesting that many experts in the fixed routing setup are underutilized, introducing redundancy as also noted in their work  (gaoHigherLayersNeed2024). In contrast, dynamic routing adapts expert selection on a token-by-token basis, which benefits complex reasoning and instruction-following tasks but may introduce variability that is less advantageous for shorter classification settings. Overall, LD-MoLE provides a stronger balance between parameter efficiency and performance, adapting better effectiveness across diverse tasks.
@@
Method Model TP MMLU_P ARC-C ARC-E Open Comm SWAG Hella CoLA RTE Avg
MoLA(8888) Llama3.1-8B 2.17 % 48.21 76.79 86.32 86.40 84.11 89.15 94.06 87.73 89.53 82.48
MoLA(2468) Llama3.1-8B 1.26 % 50.37 77.90 87.37 87.60 83.87 86.97 94.03 87.25 90.25 82.85
ReMoLE Llama3.1-8B 2.17 % 54.98 82.00 91.61 87.60 83.17 92.22 94.95 85.27 88.43 84.47
LD-MoLE Llama3.1-8B 2.33 % 55.98 83.67 91.61 88.00 83.01 92.29 95.45 85.25 91.32 85.18
MoLA(8888) Llama3.2-3B 3.11 % 40.31 71.57 83.51 81.00 79.77 83.56 87.47 85.81 90.61 78.18
MoLA(2468) Llama3.2-3B 1.80 % 42.31 71.91 83.86 83.60 80.02 83.96 87.31 86.00 89.53 78.72
ReMoLE Llama3.2-3B 3.11 % 48.01 75.25 89.30 83.40 79.52 90.45 93.44 83.95 89.46 81.42
LD-MoLE Llama3.2-3B 3.28 % 49.58 74.58 89.47 84.00 81.42 91.37 93.60 86.02 88.38 82.05
MoLA(8888) Qwen3-1.7B 4.12 % 51.12 76.59 88.60 82.40 76.49 84.11 83.35 83.89 86.64 79.24
MoLA(2468) Qwen3-1.7B 2.39 % 49.96 76.92 88.42 83.00 75.84 84.17 87.09 83.60 84.48 79.28
ReMoLE Qwen3-1.7B 4.12 % 53.40 79.60 91.75 84.80 79.44 86.37 88.00 82.12 83.74 81.02
LD-MoLE Qwen3-1.7B 4.23 % 53.82 78.67 92.11 85.00 79.30 86.72 88.71 82.61 87.72 81.63